OpenStudy (anonymous):

Fred buys several shares of stock at $24 per share and sells them at $21.12 per share. What is the percent of decrease in the price of the stock?

OpenStudy (kropot72):

Subtract the selling price from the purchase price to find the amount of decrease. Then divide the result by the purchase price and multiply that result by 100. \[\large Percentage\ decrease=\frac{24.00-21.12}{24} \times\frac{100}{1}=you\ can\ calculate\]